Once pet's age was 2 times Paul’s age. But at that time pet's age= Paul’s current age . Sum of their present ages is 40. How old is pet now?
Answer by amoresroy(361)

Once pet's age was 2 times Paul’s age. But at that time pet's age= Paul’s current age . Sum of their present ages is 40. How old is pet now?
Let x = the age of pet now
y = the age of Paul now
x-y = number of years ago when pet's age equals Paul's current age
x - (x-y) = 2[y-(x-y)]
y = 2(2y-x)
y = 4y - 2x
2x = 3y
x = 1.5y

1.5y+y = 40
y = 16
x = 24
Pet is 24 years old now.